# EvoRL: A GPU-accelerated Framework for Evolutionary Reinforcement Learning

### v2

@@ Appendix A Example Codes
EvoRL follows the object-oriented functional programming style, where classes define the running logic, and its state is externally maintained. It helps reuse components in single or batch mode. We provide an example of how the Agent and Workflow of A2C are defined and used in EvoRL.
@@
The A2CWorkflow defines the training pipeline of A2C. At its creation stage, the A2CAgent instance is created as a property of the workflow. Its step() defines one iteration of the training. For A2C, it includes: (1) sampling trajectories from environments, (2) using these trajectories as data batches to compute the loss and apply gradient-based updates for agent_state. The step() function can be JIT compiled and vectorized by JAX. In the meantime, learn() defines the training loop, which calls step() multiple times and handles the termination condition, evaluation, and stuff that cannot be JIT compiled like logging and checkpointing. The usage of A2CWorkflow also follows the object-oriented functional programming style, where its state is separately initialized and externally stored after the object instantiation.
@@ B.2. ERL Algorithms
For the original ERL and CEM-RL, we use $[256,256]$ MLP networks for actors and a shared critic. ReLU and layer normalization are applied to the hidden layers, aligning with the original implementations. The underlying RL algorithm is TD3, and no observation normalization is applied. All ERL algorithms are trained with 20000 sampled episodes. Since the RL update is applied on a shared critic, we don’t delay the actors’ update (i.e., the actor update interval is one), which is the same as the implementation in fastpbrl. Table 4 lists the hyperparameters for ERL algorithms and the underlying TD3 algorithm. To test the speed in Sec. 5.1.2, we choose fixed 4096 RL updates per iteration and record the average iteration time from 5 iterations for different implementations. The other training settings are the same as above and the time on the warm-up stage is skipped.
@@ B.3. Population-based Training
PBT Hyperparameter Value PPO Search Space Value
Population Size 128 Actor Loss Weight [0.01,10]
Warm-up Steps 256 Critic Loss Weight [0.01,10]
Workflow Steps per Iter 64 Entropy Loss Weight [-1,-1e-5]
Perturbation Factor 0.2 Discount Factor [0.86466, 0.99999]
Selection Ratio (PBT only) 0.2 GAE Factor [0.63212, 0.99999]
Episodes for Meta Obj. 16 Clip Epsilon [0.01,0.5]
Evaluation Episodes 128 Batch Size 256
Optimizer Adam
Learning Rate 3e-4
Gradient Norm. Clip 10
Timesteps per Iter 2048
Epochs per Iter 4
@@
Actor Update Interval 1

### v3

@@ 7. Conclusion
Acknowledgements.
This work was supported in part by Guangdong Basic and Applied Basic Research Foundation (No. 2024B1515020019).
@@ Appendix A Example Codes
EvoRL follows the object-oriented functional programming style, where classes define the running logic, and their state is externally maintained. It helps reuse components in single or batch mode. We provide an example of how the Agent and Workflow of A2C are defined and used in EvoRL.
@@
@classmethod
def _build_from_config(cls, config: DictConfig):
env = create_env(config.env, ...)
agent = A2CAgent(...)
optimizer = optax.adam(config.optimizer.lr)
eval_env = create_env(config.env, ...)
evaluator = Evaluator(...)
return cls(env, agent, optimizer, evaluator, config)
@@
state = state.replace(
agent_state=agent_state,
env_state=env_state,
opt_state=opt_state,
return train_metrics, state
@@
The A2CWorkflow defines the training pipeline of A2C. Its _build_from_config() defines the fundamental components in OnPolicyWorkflow, including the rollout environment, the agent object (A2CAgent) and its optimizer, the independent evaluator, and the config object during the creation stage. This method will be called when the workflow is created by its public method: A2CWorkflow.build_from_config(config). The components’ states, along with the workflow’s own states, will be initialized in init(). Its step() defines one iteration of the training. For A2C, it includes: (1) sampling trajectories from environments, (2) using these trajectories as data batches to compute the loss and apply gradient-based updates for agent_state. The step() function can be JIT compiled and vectorized by JAX. Moreover, learn() defines the training loop, which calls step() multiple times and handles the termination condition, evaluation, and stuff that cannot be JIT compiled, like logging and checkpointing. The usage of A2CWorkflow also follows the object-oriented functional programming style, where its state is separately initialized and externally stored after the object instantiation.
@@ B.2. ERL Algorithms
For the original ERL and CEM-RL, we use $[256,256]$ MLP networks for actors and a shared critic. ReLU and layer normalization are applied to the hidden layers, aligning with the original implementations. The underlying RL algorithm is TD3, and no observation normalization is applied. All ERL algorithms are trained with 20000 sampled episodes. Since the RL update is applied on a shared critic, we don’t delay the actors’ update (i.e., the actor update interval is one), which is the same as the implementation in fastpbrl. Table 5 lists the hyperparameters for ERL algorithms and the underlying TD3 algorithm. To test the speed in Sec. 5.1.2, we choose fixed 4096 RL updates per iteration and record the average iteration time from 5 iterations for different implementations. The other training settings are the same as above, and the time spent on the warm-up stage is skipped.
@@ B.3. Population-based Training
PBT Hyperparameter Value PPO Search Space Value PPO Search Space Value
Population Size 128 Actor Loss Weight [0.01,10] Batch Size 256
Warm-up Steps 256 Critic Loss Weight [0.01,10] Optimizer Adam
Workflow Steps per Iter 64 Entropy Loss Weight [-1,-1e-5] Learning Rate 3e-4
Perturbation Factor 0.2 Discount Factor [0.86466, 0.99999] Gradient Norm. Clip 10
Selection Ratio (PBT only) 0.2 GAE Factor [0.63212, 0.99999] Timesteps per Iter 2048
Episodes for Meta Obj. 16 Clip Epsilon [0.01,0.5] Number of Env. 4
Evaluation Episodes 128 Epochs per Iter 4
@@
Actor Update Interval 2
@@
Table 8. Hyperparameters for PPO
Hyperparameter Value
Actor Loss Weight 1.0
Critic Loss Weight 0.5
Entropy Loss Weight -0.01
Discount Factor 0.99
GAE Factor 0.95
Clip Epsilon 0.2
Appendix C Supplementary Experimental Results
In Sec. 5.2, we present the performance of various EvoRL algorithms on EvoRL framework, including ERL and PBT. This section provides supplementary experimental results that compare these EvoRL algorithms with their corresponding standalone RL algorithms. These comparisons further demonstrate the advantages of our framework in supporting EvoRL algorithms through end-to-end GPU acceleration.
C.1. ERL Algorithms
The ERL algorithms integrate TD3 as the underlying RL component. For comparison, we evaluate them against the original TD3 implementation under similar training settings, including identical network architectures and hyperparameters. The key differences lie in the update frequency and rollout strategies. While ERL algorithms use episodic sampling, the standalone TD3 samples two timesteps per iteration from a single environment and performs two updates (with one delayed critic update). This results in more sequential operations across the same number of episodes, thereby significantly increasing the overall training time.
Due to these distinctions in data utilization strategies, it is challenging to compare performance using a uniform horizontal axis, such as the number of sampled timesteps or episodes. Therefore, we adopt alternative perspectives to evaluate performance. Fig. 9 shows a linear interpolation of evaluation metrics during TD3 training, plotted against the number of sampled episodes to match other ERL algorithms. Compared to TD3, CEM-RL demonstrates greater training stability and achieves higher episodic returns. Although TD3 may appear to converge at a comparable or even faster rate when assessed by the number of episodes, this can be misleading, particularly in the context of end-to-end GPU acceleration and hierarchical vectorization. To address this, we conduct independent experiments using machines with identical hardware to evaluate on training time, as shown in Fig. 10. Both CEM-RL and TD3 implementations are built on EvoRL, and the reported training times exclude JIT compilation and evaluation phases. The results further highlight the superiority of CEM-RL over TD3 in terms of computational efficiency.
Figure 9. The benchmark of ERL and RL on robotic locomotion tasks. All ERL algorithms use TD3 as the RL component. Each algorithm is repeated with 8 different seeds. The average return and its 95% confidence interval about the population distribution mean are reported.
Figure 10. Evaluation of CEM-RL and TD3 on training time. Each algorithm is repeated with 4 different seeds. The average best return of the population and 95% confidence interval are reported.
C.2. PBT Algorithms
We compare the performance of PBT and PBT-CSO with the baseline PPO algorithm using fixed hyperparameters. The PPO hyperparameters, listed in Table 8, are chosen from commonly adopted values. To ensure a fair comparison under a similar level of computational resources, we train 128 PPO agents in parallel with different random seeds. This ensemble is treated as a single training instance, and we report the highest return achieved within its population. As shown in Fig. 11, both PBT and PBT-CSO significantly outperform the baseline PPO. The improvement illustrates the effectiveness of population-based hyperparameter tuning in enhancing policy performance.
Figure 11. The benchmark of PBT, PBT-CSO, and a population of PPO with fixed hyperparameters. Each algorithm is repeated with 4 different seeds. The average best return of the population and 95% confidence interval are reported.